OpenStudy (anonymous):

solve x/6 + 1/2 = 1/12

OpenStudy (anonymous):

x/6 + 1/2 = 1/12 x/6 = 1/12 - 1/2 x/6 = (2 - 12)/24 x/6 = -10/24 x = -5/2

OpenStudy (anonymous):

thank you